# Scene Graph Generation in Large-Size VHR Satellite Imagery: A Large-Scale Dataset and A Context-Aware Approach

### v2

@@ Abstract
Scene graph generation (SGG) in satellite imagery (SAI) benefits promoting intelligent understanding of geospatial scenarios from perception to cognition. In SAI, objects exhibit great variations in scales and aspect ratios, and there exist rich relationships between objects (even between spatially disjoint objects), which makes it necessary to holistically conduct SGG in large-size very-high-resolution (VHR) SAI. However, the lack of SGG datasets with large-size VHR SAI has constrained the advancement of SGG in SAI. Due to the complexity of large-size VHR SAI, mining triplets $<$subject, relationship, object$>$ in large-size VHR SAI heavily relies on long-range contextual reasoning. Consequently, SGG models designed for small-size natural imagery are not directly applicable to large-size VHR SAI. To address the scarcity of datasets, this paper constructs a large-scale dataset for SGG in large-size VHR SAI with image sizes ranging from 512 × 768 to 27,860 × 31,096 pixels, named RSG, encompassing over 210,000 objects and more than 400,000 triplets. To realize SGG in large-size VHR SAI, we propose a context-aware cascade cognition (CAC) framework to understand SAI at three levels: object detection (OBD), pair pruning and relationship prediction. As a fundamental prerequisite for SGG in large-size SAI, a holistic multi-class object detection network (HOD-Net) that can flexibly integrate multi-scale contexts is proposed. With the consideration that there exist a huge amount of object pairs in large-size SAI but only a minority of object pairs contain meaningful relationships, we design a pair proposal generation (PPG) network via adversarial reconstruction to select high-value pairs. Furthermore, a relationship prediction network with context-aware messaging (RPCM) is proposed to predict the relationship types of these pairs. To promote the development of SGG in large-size VHR SAI, this paper releases a SAI-oriented SGG toolkit with about 30 OBD methods and 10 SGG methods, and develops a benchmark based on RSG where our HOD-Net and RPCM significantly outperform the state-of-the-art methods in both OBD and SGG tasks. The RSG dataset and SAI-oriented toolkit will be made publicly available at https://linlin-dev.github.io/project/RSG.
@@ 1 Introduction
Figure 1: Illustration of SGG in large-size VHR SAI. (a) and (c) show OBD and SGG results in large-size VHR SAI, respectively. In (d), black arrows denote semantic relationships whose prediction only depends on isolated pairs, but red arrows denote semantic relationships that should be inferred with the aid of contexts.
Scene graph generation (SGG) [1, 2, 3] in satellite imagery (SAI) aims to detect objects and predict relationships between objects, where scene graph is a high-level structured representation of SAI via extensive triplets $<$subject, relationship, object$>$. Compared with traditional perceptual understanding tasks in SAI, such as object detection [4, 5, 6], scene classification [7, 8, 9] and semantic segmentation [10, 11, 12], SGG in SAI works towards cognitively mining the high-value knowledge contained in SAI (e.g., in Fig. 1, $<$ship1, away from, dock1$>$ can be inferred by the spatial layout between ship1 and dock1 and the wake trailing of ship1). SGG in SAI benefits cognitive interpretation tasks in SAI, such as image retrieval [13], image captioning [14] and image visual question answering [15], and further serves various applications, such as traffic planning [16, 17], circuit layout [18] and military situation awareness [19].
@@
As shown in Table I, existing datasets for the SGG task in SAI [20, 21] face limitations concerning the number of objects, the number of triplets, and the image size. Moreover, they tend to provide horizontal bounding box (HBB) annotations with too much background disturbance, which does not provide accurate information for recognizing relationships between man-made objects with distinct oriented boundaries in SAI. For the object detection (OBD) task in SAI, while datasets for OBD in SAI [22, 23, 5] are continually improving in object annotation granularity as well as annotation volume, they lack relationship annotations. Additionally, they also necessitate considerable improvements in terms of image size, object diversity, and scene complexity, all of which are crucial for SGG in SAI. Given these constraints, it appears unrealistic to train a robust SGG model using these limited datasets. Specifically, some special characteristics should be considered while annotating the triplets $<$subject, relationship, object$>$ for SGG in SAI: (i) Large-size VHR SAI spanning diverse scenarios. In SAI, objects exhibit large variations, and there exist rich relationships between objects in various complex scenarios, even between those not spatially intersecting. Large-size VHR SAI retains the detailed information of small objects while simultaneously preserving the integrity of large objects, which makes the predictions of relationships between objects at different scales/ranges feasible. (ii) Fine-grained objects with precise annotations. Given the top-down perspective inherent in SAI, objects such as ships, trucks and runways often appear in arbitrary orientations. Therefore, precise object annotations are necessary for accurate object localization and the exclusion of irrelevant background elements. Notably, fine-grained object categories are also indispensable for practical applications of SGG in SAI. For instance, when investigating parking violations $<$car, incorrectly parked on, truck parking$>$, it is crucial to distinguish not only between fine-grained vehicle categories (truck/car) for subject, but also between fine-grained parking categories (truck parking/car parking) for object. (iii) High-value relationships considering contextual reasoning. Unlike natural imagery, the relationships between objects in SAI are rotationally invariant, i.e., they do not change with the rotation of the imagery. In addition, many meaningful relationships between objects are heavily context-dependent, and the same types of object pairs in SAI may correspond to different categories under different contexts.
@@
TABLE I: Comparison of the RSG dataset with other datasets for OBD and SGG tasks in SAI (only for datasets sourced from satellite platforms and covering GSD $<$ 1m). The comparison includes image size, object annotation type, ground sampling distance (GSD), object information (the number of instances, the number of classes, and the number of objects per image (OPI)), relationship information (the number of triplets, the number of classes, and the number of relationships per image (RPI)), source, and public.
@@
To address the dataset scarcity problem, we construct RSG, a large-scale dataset with more than 210,000 objects and over 400,000 triplets for SGG in large-size VHR SAI. In this dataset, (i) SAI with a spatial resolution of 0.15m to 1m is collected, covering 11 categories of complex geospatial scenarios associated closely with human activities worldwide (e.g., airports, ports, nuclear power stations and dams). (ii) Under the guidance of experts in SAI, all objects are classified into 48 fine-grained categories and precisely annotated with oriented bounding boxes (OBB), and all relationships are annotated in accordance with 8 major categories including 58 fine-grained categories. (iii) All object pairs and their contained relationships are one-to-many annotated, and all relationship annotations are absolute (unaffected by imagery rotation). In conclusion, RSG has significant advantages over existing OBD datasets and SGG datasets in SAI. To the best of our knowledge, RSG is the first large-scale dataset for SGG in large-size VHR SAI.
@@
To advance the research on the cognitive understanding of SAI, this paper points out a new challenging and meaningful topic: SGG in large-size VHR SAI. Considering the characteristics of large-size VHR SAI, potential solutions should overcome three main problems: (i) OBD. Multi-class OBD in large-size VHR SAI presents significant challenges, primarily due to the holistic detection for multi-class objects with diverse object scales (e.g., same category: bridges, docks; different categories: ”airplanes and aprons”, ”ships and docks”), extreme aspect ratios (e.g., bridges, runways and gravity dams), and arbitrary orientations (e.g., trucks, ships, airplanes, bridges, runways). These factors make it challenging to holistically detect these objects while preserving the complete details necessary for accurate relationship prediction. (ii) Pair pruning. Large-size VHR SAI usually contains a large number of objects and a plethora of object pairs (the number of which grows exponentially as the number of objects), which inevitably results in excessive computation for exhaustive methods based on all object pairs. Thus, it is necessary to prune object pairs without losing high-value pairs. Notably, the triplets are usually partially annotated in reality (i.e., many non-annotated object pairs still carry meaningful relationships), which suggests that some approaches treating pair sparsity as a binary classification problem (i.e., annotated pairs are treated as positive samples and non-annotated pairs are treated as negative samples) are not reasonable. (iii) Relationship prediction. Different from small-size natural imagery, there are a large number of object pairs containing rich knowledge, which are even spatially disjoint in large-size SAI. A few relationships in large-size SAI can be directly predicted by isolated pairs, while most high-value relationships need to be inferred with the help of context. As illustrated in Fig. 1(c) and Fig. 1(d), only simple triplets (e.g., $<$crane1, over, dock1$>$) in large-size SAI can be predicted when relying only on isolated pairs. However, more complex triplets can be inferred when the context of the pairs is introduced (e.g., $<$ship2, parallelly docked at, dock1$>$ and $<$ship5, parallelly docked at, dock2$>$ can be inferred as $<$ship2, docking at the different dock with, ship5$>$).
@@
Facing the aforementioned problems, we propose a context-aware cascade cognition (CAC) framework to understand SAI at three levels: OBD, pair pruning and relationship prediction. First, a holistic multi-class object detection network (HOD-Net) incorporating multi-scale context is adopted to detect multi-scale objects in large-size VHR SAI. Then, to reduce the complexity of predicting pairs without losing high-value pairs, a pair proposal generation (PPG) network via adversarial reconstruction is designed to obtain contextual interaction pairs containing rich knowledge. Finally, considering the dependence of relationship prediction in large-size SAI on the contexts of object pairs, a relationship prediction network with context-aware messaging (RPCM) is proposed to comprehensively predict the relationship types.
@@
In summary, the RSG dataset and CAC framework proposed in this paper aim to establish a benchmark for SGG in large-size VHR SAI. With the help of this benchmark, more innovative algorithms can be developed to facilitate the cognitive understanding of SAI. The main contributions of this work are:
@@
  • We propose RSG, the first large-scale dataset for SGG in large-size VHR SAI. Containing more than 210,000 objects and over 400,000 triplets across 1,273 complex scenarios globally, the RSG dataset benefits in facilitating the development of SAI-oriented SGG.
@@
  • To address SGG in large-size VHR SAI, the CAC framework is constructed where HOD-Net is proposed for holistic detection of multi-class multi-scale objects in large-size SAI, and then the PPG network and RPCM network are designed to retain high-value pairs as well as to predict relationship types of these pairs.
@@
  • We release a versatile toolkit, encompassing about 30 OBD methods and 10 SGG methods, designed to accommodate OBD and SGG in large-size VHR SAI. The toolkit unifies the SGG process in natural imagery and SAI, offering flexibility and ease of use. Our toolkit will provide valuable contributions to the development of the SGG community.
@@
  • Rigorous experiments on the RSG dataset, employing state-of-the-art algorithms, underscore the indispensability of this dataset and the efficacy of the algorithms. The experimental results showcased in this paper serve as a benchmark for future research in this direction.
The rest of this paper is organized as follows. Section 2 reviews the related work. Section 3 introduces the proposed RSG dataset. Section 4 introduces the proposed CAC framework for SGG in large-size VHR SAI. Section 5 reports the experiments and provides a discussion of the experimental results. Section 6 gives the conclusion.
@@
Figure 2: The geographical distribution map of the sampled images from the proposed RSG dataset.
@@ 2.1 Scene Graph Generation Datasets
Scene graph has attracted extensive attention from many researchers as a powerful tool for the understanding and reasoning analysis of images [28, 29, 15, 14, 30]. Various datasets have been proposed for SGG in recent years.
2.1.1 Scene Graph Generation Datasets in Natural Imagery
@@ 2.1.2 Scene Graph Generation Datasets in SAI
Compared to the natural imagery field, the development of SGG in SAI field has been relatively slow. SGG datasets in SAI [20, 21, 40, 41] with detailed annotation and considering the characteristics of SAI are rare. GRTRD [20] may be the first SGG dataset in SAI, which focuses on 12 kinds of objects and contains 19,904 object instances and 18,602 relationship instances. RSSGD [21] is a SAI-oriented SGG dataset constructed based on the remote sensing caption dataset RSICD [42], which has about 39 object categories and 16 relationship categories. Unfortunately, the image size of GRTRD dataset is 600×600 pixels, and the image size of RSSGD dataset is only 224×224 pixels. Since the large-size VHR SAI scenarios are cropped into small image blocks, the objects and relationships contained in each image are very sparse, and many object pairs containing high-value relationships are corrupted, which makes it difficult to support high-level cognitive understanding of SAI. In addition, the objects in the above dataset are annotated by HBB with too much background disturbance, which does not provide accurate information for recognizing relationships between regular man-made objects with distinct oriented boundaries in large-size VHR SAI. To meet the need for cognitive understanding of SAI, a large-scale SGG dataset named RSG for large-size VHR SAI is constructed, which contains multiple complex scenarios and meaningful relationships. In Table I, we give the statistics of popular OBD datasets and SGG datasets in SAI. Compared to small-scale datasets for the SGG task in SAI, such as RSSGD dataset and GRTRD dataset, our RSG outperforms them by at least an order of magnitude in both the number of objects and relationships on all/single images. For OBD datasets in SAI [24, 25, 23] lacking relationship annotations, our RSG dataset provides effective improvements in both object category diversity and scene complexity. The RSG dataset, a large-scale and comprehensive SGG dataset, supports both the OBD task and the SGG task based on the HBB-based and OBB-based detectors in large-size VHR SAI, serving as an invaluable resource for cognitive understanding of SAI.
@@ 2.2 Scene Graph Generation Methods
Figure 3: Statistics and visualization of objects (a) and relationships (b) from the RSG dataset. The relationships are color-coded to show parking statu, spatial topology and functional description, movement status, distance warning, circuit layout, construction status and emission status. Some typical objects and interesting triplets are visualized.

### v3

# STAR: A First-Ever Dataset and A Large-Scale Benchmark for Scene Graph Generation in Large-Size Satellite Imagery

@@ Abstract
Scene graph generation (SGG) in satellite imagery (SAI) benefits promoting understanding of geospatial scenarios from perception to cognition. In SAI, objects exhibit great variations in scales and aspect ratios, and there exist rich relationships between objects (even between spatially disjoint objects), which makes it attractive to holistically conduct SGG in large-size very-high-resolution (VHR) SAI. However, there lack such SGG datasets. Due to the complexity of large-size SAI, mining triplets $<$subject, relationship, object$>$ heavily relies on long-range contextual reasoning. Consequently, SGG models designed for small-size natural imagery are not directly applicable to large-size SAI. This paper constructs a large-scale dataset for SGG in large-size VHR SAI with image sizes ranging from 512 × 768 to 27,860 × 31,096 pixels, named STAR (Scene graph generaTion in lArge-size satellite imageRy), encompassing over 210K objects and over 400K triplets. To realize SGG in large-size SAI, we propose a context-aware cascade cognition (CAC) framework to understand SAI regarding object detection (OBD), pair pruning and relationship prediction for SGG. We also release a SAI-oriented SGG toolkit with about 30 OBD and 10 SGG methods which need further adaptation by our devised modules on our challenging STAR dataset. The dataset and toolkit are available at: https://linlin-dev.github.io/project/STAR.
@@ 1 Introduction
Figure 1: Illustration of scene graph generation (SGG) in large-size VHR satellite imagery (SAI). (a) and (c) show OBD and SGG results in large-size VHR SAI, respectively. In (d), black arrows denote semantic relationships whose prediction only depends on isolated pairs, but red arrows denote semantic relationships that as inferred from contexts.
Scene graph generation (SGG) [1, 2, 3] in satellite imagery (SAI) aims to detect objects and predict relationships between objects, where scene graph is a high-level structured representation of SAI via extensive triplets $<$subject, relationship, object$>$. Compared with traditional perceptual understanding tasks in SAI, such as object detection [4, 5, 6], scene classification [7, 8, 9] and semantic segmentation [10, 11, 12], SGG in SAI works towards cognitively mining the high-value knowledge contained in SAI (e.g., in Fig. 1, $<$ship1, away from, dock1$>$ can be inferred by the spatial layout between ship1 and dock1 and the wake trailing of ship1). SGG in SAI benefits cognitive interpretation tasks in SAI, like image retrieval [13], image captioning [14] and image visual question answering [15], and further serves various applications, e.g.traffic planning [16, 17], circuit layout [18].
@@
Table I shows existing datasets for the SGG task in SAI [19, 20], which face limitations concerning the number of objects, number of triplets, and image size. Moreover, they tend to provide horizontal bounding box (HBB) annotations with too much background disturbance, hurting the recognition of relationships between man-made objects with distinct oriented boundaries in SAI. For the object detection (OBD) task in SAI, while the datasets for OBD in SAI [21, 22, 5] are continually improving in object annotation granularity as well as annotation volume, they lack relationship annotations. Additionally, they necessitate considerable improvements in terms of image size, object diversity, and scene complexity, all of which are crucial for the SGG task in SAI. Given these constraints, it appears unrealistic to train a robust SGG model using these limited datasets. Specifically, some specific characteristics should be considered while annotating the triplets $<$subject, relationship, object$>$ for the SGG task in SAI: (i) Large-size VHR SAI spanning diverse scenarios. In SAI, objects exhibit large variations, and there exist rich relationships between objects in various complex scenarios, even between those not spatially intersecting. Large-size VHR SAI retains the detailed information of small objects while simultaneously preserving the integrity of large objects, which makes the predictions of relationships between objects at different scales/ranges feasible. (ii) Fine-grained objects with precise annotations. Given the top-down perspective inherent in SAI, objects such as ships, trucks and runways often appear in arbitrary orientations. Therefore, precise object annotations are necessary for accurate localization. Notably, fine-grained object categories are also indispensable for practical applications of SGG task in SAI. For instance, when investigating parking violations $<$car, incorrectly parked on, truck parking$>$, it is crucial to distinguish not only between fine-grained vehicle categories (truck/car) for subject, but also between fine-grained parking categories (truck parking/car parking) for object. (iii) High-value relationships considering contextual reasoning. Unlike natural imagery, the relationships between objects in SAI are rotationally invariant (i.e., they do not change with the rotation of the image), and a significant number of object pairs containing high-value relationships are spatially disjoint in SAI. Furthermore, many meaningful relationships between objects are heavily context-dependent, and the same types of object pairs in SAI may correspond to different categories under different contexts.
@@
TABLE I: Comparison of the STAR dataset with other datasets for OBD and SGG tasks in SAI (only for datasets sourced from satellite platforms and covering GSD $<$ 1m). The comparison includes image size, object annotation type, ground sampling distance (GSD), object information (the number of instances, the number of classes, and the number of objects per image (OPI)), relationship information (the number of triplets, the number of classes, and the number of relationships per image (RPI)), source, and public.
@@
To address the dataset scarcity, we introduce a first-ever dataset for Scene graph generaTion in lArge-size satellite imageRy (STAR), which contains more than 210,000 objects and more than 400,000 triplets for SGG in large-size VHR SAI. In this dataset, (i) SAI with a spatial resolution of 0.15m to 1m is collected, covering 11 categories of complex geospatial scenarios associated closely with human activities worldwide (e.g., airports, ports, nuclear power stations and dams). (ii) By the guidance of human experts, all objects are classified into 48 fine-grained categories and annotated with oriented bounding boxes (OBB), and all relationships are annotated in accordance with 8 major categories including 58 fine-grained categories. (iii) All object pairs and their contained relationships are one-to-many annotated, and all relationship annotations are absolute (unaffected by imagery rotation). In conclusion, STAR has significant advantages over existing OBD datasets and SGG datasets in SAI.
@@
Specifically, this paper opens up a new challenging and practical topic: SGG in large-size VHR SAI. Considering the characteristics of large-size VHR SAI, potential solutions should overcome three main problems: (i) OBD. Multi-class OBD in large-size VHR SAI presents significant challenges, primarily due to the holistic detection for multi-class objects with diverse object scales (e.g., same category: bridges, docks; different categories: ”airplanes and aprons”, ”ships and docks”), extreme aspect ratios (e.g., bridges, runways and gravity dams), and arbitrary orientations (e.g., trucks, ships, airplanes, bridges, runways). These factors make it challenging to holistically detect these objects while preserving the complete details necessary for accurate relationship prediction. (ii) Pair pruning. Large-size VHR SAI usually contains a large number of objects and a plethora of object pairs (the number of which grows exponentially as the number of objects), which inevitably results in excessive computation for exhaustive methods based on all object pairs. Thus, it is necessary to prune object pairs without losing high-value pairs. Notably, the triplets are usually partially annotated in reality (i.e., many non-annotated object pairs still carry meaningful relationships), which suggests that some approaches treating pair sparsity as a binary classification problem (i.e., annotated pairs are treated as positive samples and non-annotated pairs are treated as negative samples) are not reasonable. (iii) Relationship prediction. Different from small-size natural imagery, there are a large number of object pairs containing rich knowledge, which are even spatially disjoint in large-size SAI. A few relationships in large-size SAI can be directly predicted by isolated pairs, while most high-value relationships need to be inferred with the help of context. As illustrated in Fig. 1(c) and Fig. 1(d), only simple triplets (e.g., $<$crane1, over, dock1$>$) in large-size SAI can be predicted when relying only on isolated pairs. However, more complex triplets can be inferred when the context of the pairs is introduced (e.g., $<$ship2, parallelly docked at, dock1$>$ and $<$ship5, parallelly docked at, dock2$>$ can be inferred as $<$ship2, docking at the different dock with, ship5$>$).
@@
In this paper, we propose a context-aware cascade cognition (CAC) framework to understand SAI at three levels: OBD, pair pruning and relationship prediction. First, a holistic multi-class object detection network (HOD-Net) incorporating multi-scale context is adopted to detect multi-scale objects in large-size VHR SAI. Then, to reduce the complexity of predicting pairs without losing high-value pairs, a pair proposal generation (PPG) network via adversarial reconstruction is designed to obtain contextual interaction pairs containing rich knowledge. Finally, considering the dependence of relationship prediction in large-size SAI on the contexts of object pairs, a relationship prediction network with context-aware messaging (RPCM) is proposed to comprehensively predict the relationship types.
@@
In summary, the STAR dataset and CAC framework proposed in this paper aim to establish a large-scale benchmark for SGG in large-size VHR SAI. Under this benchmark, new algorithms can be further developed. The contributions are:
@@
  • We propose STAR, a first-ever dataset for SGG in large-size VHR SAI. Containing more than 210K objects and over 400K triplets across 1,273 complex scenarios globally, it provides more challenging and practical testbed for the community.
@@
  • To address SGG in large-size VHR SAI, a context-aware cascade cognition (CAC) framework is established where the HOD-Net is devised for holistic detection of multi-class multi-scale objects in large-size SAI, and then the pair proposal generation (PPG) network and relationship prediction network with context-aware messaging (RPCM) network are designed to retain high-value pairs as well as to predict relationship types of these pairs.
@@
  • We release an open-source toolkit, encompassing about 30 OBD methods and 10 SGG methods, designed to accommodate OBD and SGG in large-size VHR SAI. The versatile toolkit unifies the SGG process in natural imagery and SAI, offering flexibility and ease of use. Our toolkit will provide valuable contributions to the development of the SGG community.
@@
  • Rigorous experiments on the STAR dataset, employing state-of-the-art algorithms, underscore the indispensability of this dataset and the efficacy of the algorithms. The experimental results showcased in this paper serve as a large-scale benchmark for future research in this direction.
@@
Figure 2: The geographical distribution map of the sampled images from the proposed STAR dataset.
@@ 2.1 Scene Graph Generation Datasets
Scene graph has attracted extensive attention from many researchers as a powerful tool for the understanding and reasoning analysis of images [27, 28, 15, 14]. Various datasets have been proposed for SGG in recent years.
2.1.1 Scene Graph Generation Datasets in Natural Images
@@ 2.1.2 Scene Graph Generation Datasets in SAI
Compared to the natural imagery field, the development of SGG in SAI field has been relatively slow. SGG datasets in SAI [19, 20, 38, 39] with detailed annotation and considering the characteristics of SAI are rare. GRTRD [19] may be the first SGG dataset in SAI, which focuses on 12 kinds of objects and contains 19,904 object instances and 18,602 relationship instances. RSSGD [20] is a SAI-oriented SGG dataset constructed based on the remote sensing caption dataset RSICD [40], which has about 39 object categories and 16 relationship categories. Unfortunately, the image size of GRTRD dataset is 600×600 pixels, and the image size of RSSGD dataset is only 224×224 pixels. Since the large-size VHR SAI scenarios are cropped into small image blocks, the objects and relationships contained in each image are very sparse, and many object pairs containing high-value relationships are corrupted, which makes it difficult to support high-level cognitive understanding of SAI. In addition, the objects in the above dataset are annotated by HBB with too much background disturbance, which does not provide accurate information for recognizing relationships between regular man-made objects with distinct oriented boundaries in large-size VHR SAI. To meet the need for cognitive understanding of SAI, a large-scale SGG dataset named STAR for large-size VHR SAI is constructed, which contains multiple complex scenarios and meaningful relationships. In Table I, we give the statistics of popular OBD datasets and SGG datasets in SAI. Compared to small-scale datasets for the SGG task in SAI, such as RSSGD dataset and GRTRD dataset, our STAR outperforms them by at least an order of magnitude in both the number of objects and relationships on all/single images. For OBD datasets in SAI [23, 24, 22] lacking relationship annotations, our STAR dataset provides effective improvements in both object category diversity and scene complexity. The STAR dataset, a large-scale and comprehensive SGG dataset, supports both the OBD task and the SGG task based on the HBB-based and OBB-based detectors in large-size VHR SAI, serving as an invaluable resource for cognitive understanding of SAI.
@@ 2.2 Scene Graph Generation Methods
Figure 3: Statistics and visualization of objects (a) and relationships (b) from the STAR dataset. The relationships are color-coded to show parking statu, spatial topology and functional description, movement status, distance warning, circuit layout, construction status and emission status. Some typical objects and interesting triplets are visualized.